# Robust covariance estimation and explainable outlier detection for matrix-valued data

### v1

@@ 3 Properties of the MMCD estimators
As outlined in Lopuhaa and Rousseeuw, (1991), the MCD estimator has the highest possible breakdown point an affine equivariant estimator can attain. Now that we have generalized the MCD estimator to the matrix-variate setting, we want to study how the breakdown point and equivariance properties translate to this setting.
Affine equivariance in the multivariate setting can be motivated by the fact that location and covariance estimators should transform in the same way under affine transformations as the parameters of elliptically symmetrical unimodal distributions (from here on called elliptical distributions), see Maronna et al., (2019). We can define the matrix-variate analog of affine equivariance based on the properties of elliptical distributions, which are frequently employed to study robustness properties of normal theory under nonnormal situations (Gupta and Nagar,, 1999).
Using the same notation and parameters as for the matrix normal distribution, a random matrix $\bm{X}$ with an absolutely continuous distribution has an elliptical distribution, denoted $\mathcal{ME}(\bm{\operatorname{M}},\bm{\Sigma}^{\operatorname{row}},\bm{\Sigma}^{\operatorname{col}})$, if its density can be written as
$f(\bm{X})=\det(\bm{\Sigma}^{\operatorname{row}})^{\nicefrac{{q}}{{2}}}\det(\bm{\Sigma}^{\operatorname{col}})^{\nicefrac{{p}}{{2}}}g(\operatorname{tr}(\bm{\Omega}^{\operatorname{col}}(\bm{X}-\bm{\operatorname{M}})^{\prime}\bm{\Omega}^{\operatorname{row}}(\bm{X}-\bm{\operatorname{M}}))),$ (15)
with any positive $g$ such that $f(\bm{X})$ integrates to one. Linear functions of a random matrix $\bm{X}\sim\mathcal{ME}(\bm{\operatorname{M}},\bm{\Sigma}^{\operatorname{row}},\bm{\Sigma}^{\operatorname{col}})$ also have an elliptical distribution (Gupta and Varga,, 2012). This means that for constant matrices $\bm{A}\in\mathbb{R}^{r\times p}$, $\operatorname{rank}(\bm{A})=r\leq p$, $\bm{B}\in\mathbb{R}^{q\times s}$, $\operatorname{rank}(\bm{B})=s\leq q$, and $\bm{C}\in\mathbb{R}^{r\times s}$, the transformed random matrix $\bm{Z}=\bm{A}\bm{X}\bm{B}+\bm{C}$ has density
$\bm{Z}\sim E_{r,s}(\bm{A}\bm{\operatorname{M}}\bm{B}+\bm{C},\bm{A}\bm{\Sigma}^{\operatorname{row}}\bm{A}^{\prime},\bm{B}^{\prime}\bm{\Sigma}^{\operatorname{col}}\bm{B}).$ (16)
@@
Properties (17) provide a suitable generalization of affine equivariance to the matrix-variate setting, and it is easy to verify that they hold for the MLEs. However, they do not imply affine equivariance of the location and covariance estimators for the vectorized observations. This would only hold for transformations with the Kronecker structure $\operatorname{vec}(\bm{A}\bm{X}\bm{B}+\bm{C})=(\bm{B}^{\prime}\otimes\bm{A})\operatorname{vec}(\bm{X})+\operatorname{vec}(\bm{C})$. To avoid confounding definitions, we refer to Properties (17) as matrix affine equivariance.
@@
Lemma 3.0.1 shows that the MMCD estimators are equivariant under linear matrix transformations, and a proof is given in Supplement B.1.
@@
A proof building on similar methodologies established in Rousseeuw, (1985) and Lopuhaa and Rousseeuw, (1991) is given in Supplement B.1. Theorem 3.0.1 implies that the maximum breakdown point of the MMCD estimators is $\nicefrac{{1}}{{n}}\lfloor\nicefrac{{(n-d)}}{{2}}\rfloor$ and is attained if $h=\lfloor\nicefrac{{(n+d+2)}}{{2}}\rfloor$. This implies that the maximum breakdown point of the MMCD covariance estimators for $p\geq 2,q\geq 2$ is higher than the upper bound for the breakdown point of affine equivariant covariance estimators applied to vectorized samples, which is given by $\nicefrac{{1}}{{n}}\lfloor\nicefrac{{(n-pq+1)}}{{2}}\rfloor$ (Davies,, 1987; Lopuhaa and Rousseeuw,, 1991). However, as mentioned earlier, affine equivariance in the matrix-variate setting does not imply affine equivariance in the multivariate setting. Thus, the mentioned upper bound for the vectorized observations does not apply. In other words, since affine equivariance (in the vectorized case) is not a requirement for matrix-variate affine equivariance, it is possible to achieve a higher breakdown point for the MMCD estimators than for the MCD estimators of vectorized samples.
@@
By multiplying the raw MMCD rowwise or columnwise covariance estimator by a consistency factor $c(\alpha)$, we can obtain consistency at the normal model as outlined in Theorem 3.0.2. The proof of this theorem shows that the consistency factor of the MCD estimator and the MMCD estimator must coincide, and therefore, we use the consistency factor
$c(\alpha)=\frac{\alpha}{F_{\chi^{2}_{p+2}}(\chi^{2}_{\alpha;pq})}$ (20)
proposed by Croux and Haesbroeck, (1999), where $F_{\chi^{2}_{p+2}}$ denotes the CDF of the chi-square distribution with $p+2$ degrees of freedom, and $\chi^{2}_{pq;\alpha}$ denotes the $\alpha$ quantile of the chi-square distribution with $pq$ degrees of freedom.
@@
Let $\bm{X}_{1},\dots,\bm{X}_{n}$ be a random sample from an elliptical matrix-variate distribution $\mathcal{ME}(\bm{\operatorname{M}},\bm{\Sigma}^{\operatorname{row}},\bm{\Sigma}^{\operatorname{row}})$ with positive definite covariances $\bm{\Sigma}^{\operatorname{row}},\bm{\Sigma}^{\operatorname{col}}$, and let $(\hat{\bm{\operatorname{M}}},\hat{\bm{\Sigma}}^{\operatorname{row}},\hat{\bm{\Sigma}}^{\operatorname{col}})$ be the corresponding MMCD estimators. Then it holds that
$\|\hat{\bm{\operatorname{M}}}-\bm{\operatorname{M}}\|\xrightarrow{a.s.}0,\quad\left|c_{1}\hat{\bm{\Sigma}}^{\operatorname{col}}\otimes\hat{\bm{\Sigma}}^{\operatorname{row}}-\bm{\Sigma}^{\operatorname{col}}\otimes\bm{\Sigma}^{\operatorname{row}}\right|\xrightarrow{a.s.}0,$
where $c_{1}$ is a distribution-specific consistency factor as in Croux and Haesbroeck, (1999).
The proof of the consistency of the raw MMCD estimators relies on the strong consistency of the MCD estimator given in Butler et al., (1993); Cator and Lopuhaä, (2012) and is provided in Supplement B.1.
In the multivariate case, Lopuhaa and Rousseeuw, (1991) proposed a one-step reweighting approach to improve the efficiency of the MCD estimator and showed that the breakdown point of the original estimator is preserved. We generalize this approach to the matrix-variate setting to obtain the reweighted MMCD estimator.
@@
$\tilde{\bm{\operatorname{M}}}$ $=\frac{1}{\sum_{i=1}^{n}w(d(\bm{X}_{i}))}\sum_{i=1}^{n}w(d(\bm{X}_{i}))\bm{X}_{i},$ (21)
$\tilde{\bm{\Sigma}}^{\operatorname{row}}$ $=\frac{1}{q\sum_{i=1}^{n}w(d(\bm{X}_{i}))}\sum_{i=1}^{n}w(d(\bm{X}_{i}))(\bm{X}_{i}-\tilde{\bm{\operatorname{M}}})\tilde{\bm{\Omega}}^{\operatorname{col}}(\bm{X}_{i}-\tilde{\bm{\operatorname{M}}})^{\prime}\text{, and}$ (22)
$\tilde{\bm{\Sigma}}^{\operatorname{col}}$ $=\frac{1}{p\sum_{i=1}^{n}w(d(\bm{X}_{i}))}\sum_{i=1}^{n}w(d(\bm{X}_{i}))(\bm{X}_{i}-\tilde{\bm{\operatorname{M}}})^{\prime}\tilde{\bm{\Omega}}^{\operatorname{row}}(\bm{X}_{i}-\tilde{\bm{\operatorname{M}}}),$ (23)
where $w:[0,\infty)\rightarrow[0,\infty)$ is a non-increasing and bounded weight function such that $w(d(\bm{X}_{i}))>0$ for at least $\lfloor\nicefrac{{(n+d+2)}}{{2}}\rfloor$ observations that vanishes for large distances, i.e., $w(d(\bm{X}_{i}))=0$ if $d(\bm{X}_{i})>c_{1}>0$.
@@
[…]
@@ 4.1 Adapting the C-step
Adapting the structure of the Fast-MCD algorithm to the matrix-variate setting leads to the development of the Fast-MMCD algorithm. This adaptation necessitates a modification in the covariance estimation during the C-step to derive suitable counterparts for computing the MMCD estimators. However, this process encounters a challenge due to the involvement of two covariance matrices, as depicted in Equations (10) and (11), both lacking closed-form solutions for their estimation. To address this issue, we incorporate the flip-flop algorithm introduced by Dutilleul, (1999) within the C-step. Consider a matrix-variate random sample $\bm{\mathfrak{X}}=(\bm{X}_{1},\dots,\bm{X}_{n})$, with $\bm{X}_{i}\in\mathbb{R}^{p\times q}$, and any $h$-subset $H_{\operatorname{old}}\subset\{1,\ldots,n\}$, with $\left|H_{\operatorname{old}}\right|=h>\lfloor\nicefrac{{p}}{{q}}+\nicefrac{{q}}{{p}}\rfloor+2$. First, the MLEs $(\hat{\bm{\operatorname{M}}}_{H_{\operatorname{old}}},\hat{\bm{\Sigma}}^{\operatorname{row}}_{H_{\operatorname{old}}},\hat{\bm{\Sigma}}^{\operatorname{col}}_{H_{\operatorname{old}}})$ are computed based on the observations in the subset $H_{\operatorname{old}}$ using the flip-flop algorithm, which is non-decreasing in likelihood. Next, compute the squared Mahalanobis distances $d^{2}_{i}(H_{\operatorname{old}}):=\operatorname{MMD}^{2}(\bm{X}_{i},\hat{\bm{\operatorname{M}}}_{H_{\operatorname{old}}},\hat{\bm{\Sigma}}^{\operatorname{row}}_{H_{\operatorname{old}}},\hat{\bm{\Sigma}}^{\operatorname{col}}_{H_{\operatorname{old}}})$ for all $i=1,\ldots,n$. In Proposition 2.0.1, we showed that $\sum_{i\in H_{\operatorname{old}}}d^{2}_{i}(H_{\operatorname{old}})=hpq$, hence only the terms of the log-likelihood function involving the determinants change in this step. To construct the new subset $H_{\operatorname{new}}$, sort the squared MMDs in ascending order, resulting in a permutation $\pi$ of $\{1,\ldots,n\}$ such that $d^{2}_{\pi(1)}(H_{\operatorname{old}})\leq\ldots\leq d^{2}_{\pi(n)}(H_{\operatorname{old}})$, and define a new $h$-subset $H_{\operatorname{new}}=\{\pi(1),\ldots,\pi(h)\}$. Since the estimators do not change in this step, the terms involving the determinant are constant and by construction, the sum of the Mahalanobis distances either decreases or stays constant. Hence, the reordering is non-decreasing in likelihood. Finally, the MLEs are updated using the flip-flop algorithm based on the observations in the subset $H_{\operatorname{new}}$, resulting in estimators $(\hat{\bm{\operatorname{M}}}_{H_{\operatorname{new}}},\hat{\bm{\Sigma}}^{\operatorname{row}}_{H_{\operatorname{new}}},\hat{\bm{\Sigma}}^{\operatorname{col}}_{H_{\operatorname{new}}})$, increasing the likelihood once more, and it follows that
@@ 4.2 The Fast-MMCD algorithm
The Fast-MMCD algorithm is a matrix-variate extension of the Fast-MCD procedure of Rousseeuw and Driessen, (1999), aiming to alleviate the C-steps dependence on the initial subset by using multiple initial subsets, iteratively conducting C-steps on each until convergence, and ultimately selecting the solution with the lowest determinant. While this explains the idea of the algorithm, there are more computational considerations and adjustments in the full Fast-MMCD. A pseudo-code of the FastMMCD Algorithm 2 is given in Supplement C.
@@
As in its multivariate counterpart, the Fast-MMCD procedure uses so-called elemental subsets to initialize the procedure. This means that we use $m$ subsets of size $d+2,d=\lfloor\nicefrac{{p}}{{q}}+\nicefrac{{q}}{{p}}\rfloor$, instead of size $h$, to increase the probability of obtaining at least one clean initial subset. Using $m=500$ elemental subsets by default allows for a reasonable tradeoff between a wide variety of settings where we likely obtain at least one clean elemental subset and the computational demands of computing initial estimators. If either $p\ll q$ or $q\gg p$, $d$ will be large, and using more initial subsets is recommended. Using elemental subsets increases not only the robustness of the initial estimators but also the computational efficiency.
@@
Moreover, the Fast-MMCD procedure only uses 2 C-step and MLE iterations for the initial elemental subsets to ensure even faster computation of the initial estimators. In the MLE procedure, Werner et al., (2008) demonstrated that the same asymptotic efficiency can be attained using only two iterations instead of iterating until convergence. As for the C-step, Rousseeuw and Driessen, (1999) outlined that after two iterations, subsets with the lowest covariance determinant during the procedure can already be identified, even before reaching convergence. Moreover, simulations show that we can identify those initial subsets that yield robust solutions after 2 C-steps, whether we use 2 MLE iterations or iterate the flip-flop algorithm until convergence. This is described in detail in Supplement C.1, where we also show that elemental subsets indeed yield more robust solutions than their larger counterparts in case of high contamination.
@@
The initialization step of the FastMMCD procedure yields $m$ initial estimators and we keep the $10$ estimators with the lowest covariance determinant. Using those as initial estimators, we iterate C-steps until convergence on the full data set $\bm{\mathfrak{X}}$. The solution with the lowest covariance determinant then yields the raw MMCD estimators.
@@
The raw MMCD estimators are scaled using the consistency factor $c(\alpha)$ given in Equation (20) to achieve consistency at the normal model as outlined in Theorem 3.0.2. Based on those rescaled raw MMCD estimators, the reweighted estimators described in Definition 3.0.1 are computed using the weights given in Equation (24). The reweighted MMCD estimators are then scaled using $c(\tilde{\alpha})=c(\nicefrac{{\tilde{h}}}{{n}})$, where $\tilde{h}$ denotes the number of observations with weights one. Simulations in Section 6 show that the finite sample efficiency of the reweighed MMCD estimators increases to over $90\%$ as the sample size increases.
@@
The Fast-MMCD algorithm repeatedly computes Mahalanbois distances for all $n$ samples, which is computationally expensive when $n$ gets large. To improve the computational efficiency for settings where $n$ is large, we implemented the subsampling approach proposed by Rousseeuw and Driessen, (1999). The idea is to split the sample of $n$ observations into several smaller subsamples and compute initial estimators on those subsamples before working on the large set with $n$ observations.

### v2

@@ 3 Properties of the MMCD estimators
Matrix affine equivariance.
The concept of affine equivariance in multivariate analysis is rooted in the idea that the estimators used for location and covariance should transform in the same way as the parameters of elliptically symmetrical unimodal distributions (referred to as elliptical distributions hereafter), see Maronna et al., (2019). We can define the matrix-variate analog of affine equivariance based on the properties of matrix-variate elliptical distributions, which are frequently employed to study the robustness properties of normal theory under nonnormal situations (Gupta and Nagar,, 1999).
Linear functions of a random matrix $\bm{X}\sim\mathcal{ME}(\bm{\operatorname{M}},\bm{\Sigma}^{\operatorname{row}},\bm{\Sigma}^{\operatorname{col}},g)$ also have an elliptical distribution (Gupta and Varga,, 2012). This means that for constant matrices $\bm{A}\in\mathbb{R}^{r\times p}$, $\operatorname{rank}(\bm{A})=r\leq p$, $\bm{B}\in\mathbb{R}^{q\times s}$, $\operatorname{rank}(\bm{B})=s\leq q$, and $\bm{C}\in\mathbb{R}^{r\times s}$, the transformed random matrix $\bm{Z}=\bm{A}\bm{X}\bm{B}+\bm{C}$ has density
$\bm{Z}\sim\mathcal{ME}(\bm{A}\bm{\operatorname{M}}\bm{B}+\bm{C},\bm{A}\bm{\Sigma}^{\operatorname{row}}\bm{A}^{\prime},\bm{B}^{\prime}\bm{\Sigma}^{\operatorname{col}}\bm{B},g).$ (14)
@@
Properties (15) provide a suitable generalization of affine equivariance to the matrix-variate setting, and it is easy to verify that they hold for the estimators given in (3)-(5). However, they do not imply affine equivariance of the location and covariance estimators for the vectorized observations. This would only hold for transformations with the Kronecker structure $\operatorname{vec}(\bm{A}\bm{X}\bm{B}+\bm{C})=(\bm{B}^{\prime}\otimes\bm{A})\operatorname{vec}(\bm{X})+\operatorname{vec}(\bm{C})$. We refer to Properties (15) as matrix affine equivariance to avoid confounding definitions.
@@
Lemma 3.0.1 shows that the MMCD estimators are equivariant under matrix affine transformations, and a proof is given in Supplement B.1.
Breakdown point.
@@
The proof extends established methodologies from Rousseeuw, (1985) and Lopuhaa and Rousseeuw, (1991) to address the matrix variate setting, leveraging additional insights and techniques outlined in Supplement B.1. Theorem 3.0.1 implies that the maximum breakdown point of the MMCD estimators is $\nicefrac{{1}}{{n}}\lfloor\nicefrac{{(n-d)}}{{2}}\rfloor$ and is attained if $h=\lfloor\nicefrac{{(n+d+2)}}{{2}}\rfloor$. This means that the maximum breakdown point of the MMCD covariance estimators for $p\geq 2,q\geq 2$ is higher than the upper bound for the breakdown point of affine equivariant covariance estimators applied to vectorized samples, which is given by $\nicefrac{{1}}{{n}}\lfloor\nicefrac{{(n-pq+1)}}{{2}}\rfloor$ (Davies,, 1987; Lopuhaa and Rousseeuw,, 1991). However, as mentioned earlier, affine equivariance in the matrix-variate setting does not imply affine equivariance in the multivariate setting. Thus, the mentioned upper bound for the vectorized observations does not apply. In other words, since affine equivariance (in the vectorized case) is not a requirement for matrix-variate affine equivariance, it is possible to achieve a higher breakdown point for the MMCD estimators than for any affine equivariant multivariate estimator applied to the vectorized data.
@@
Consistency for elliptical distributions.
Let us now consider the asymptotic behavior of the MMCD estimators. By scaling the rowwise or columnwise MMCD covariance estimator by a distribution-specific consistency factor, we can achieve consistency for elliptical distributions.
@@
Let $\bm{X}_{1},\dots,\bm{X}_{n}$ be a random sample from an elliptical matrix-variate distribution $\mathcal{ME}(\bm{\operatorname{M}},\bm{\Sigma}^{\operatorname{row}},\bm{\Sigma}^{\operatorname{row}},g)$ with positive definite covariances $\bm{\Sigma}^{\operatorname{row}},\bm{\Sigma}^{\operatorname{col}}$, and let $(\hat{\bm{\operatorname{M}}},\hat{\bm{\Sigma}}^{\operatorname{row}},\hat{\bm{\Sigma}}^{\operatorname{col}})$ be the corresponding MMCD estimators. Then, it holds that
$\left\|\hat{\bm{\operatorname{M}}}-\bm{\operatorname{M}}\right\|\xrightarrow{a.s.}0,\quad\left\|{c(\alpha)\hat{\bm{\Sigma}}^{\operatorname{col}}\otimes\hat{\bm{\Sigma}}^{\operatorname{row}}-\bm{\Sigma}^{\operatorname{col}}\otimes\bm{\Sigma}^{\operatorname{row}}}\right\|\xrightarrow{a.s.}0,$
where $c(\alpha),\alpha=\nicefrac{{h}}{{n}}\in[0.5,1],$ is a distribution-specific consistency factor as in Croux and Haesbroeck, (1999).
The proof of the consistency of the raw MMCD estimators relies on the strong consistency of the MCD estimator given in Butler et al., (1993); Cator and Lopuhaä, (2012) and is provided in Supplement B.1. It shows that the consistency factor of the MCD estimator and the MMCD estimator must coincide, and therefore, we use the consistency factor
$c(\alpha)=\frac{\alpha}{F_{\chi^{2}_{pq+2}}(\chi^{2}_{\alpha;pq})}$ (18)
proposed by Croux and Haesbroeck, (1999) to obtain consistency at the normal model, where $F_{\chi^{2}_{pq+2}}$ denotes the CDF of the chi-square distribution with $pq+2$ degrees of freedom, and $\chi^{2}_{pq;\alpha}$ denotes the $\alpha$ quantile of the chi-square distribution with $pq$ degrees of freedom.
Remark 3.0.1.
Note first that for $h=n$, the corresponding MMCD estimators coincide with the ones defined in (3)-(5). Therefore, a simple, yet not discussed in the literature, consequence of Theorem 3.0.2 is that the estimators obtained maximizing the likelihood under the matrix-normal model, are consistent estimators of the corresponding finite-dimensional parameters $(\bm{\operatorname{M}},\bm{\Sigma}^{\operatorname{row}},\bm{\Sigma}^{\operatorname{row}})$ in the semi-parametric, matrix elliptical family.
Reweighted MMCD - improving efficiency.
The raw MMCD estimators are most robust when about half of the observations are trimmed, i.e., $h=\lfloor\nicefrac{{(n+d+2)}}{{2}}\rfloor$. However, this leads to a low efficiency at the normal model. While efficiency could be increased by trimming fewer samples, this would lead to lower robustness. To enhance a robust estimator’s efficiency without compromising robustness, Lopuhaa and Rousseeuw, (1991); Maronna et al., (2019) proposed a one-step reweighing procedure. We can apply this technique for the MMCD estimators by defining weighted ML estimators with weights depending on the Mahalanobis distances given the raw MMCD estimators.
@@
$\tilde{\bm{\operatorname{M}}}$ $=\frac{1}{\sum_{i=1}^{n}w(\operatorname{MMD}(\bm{X}_{i}))}\sum_{i=1}^{n}w(\operatorname{MMD}(\bm{X}_{i}))\bm{X}_{i},$ (19)
$\tilde{\bm{\Sigma}}^{\operatorname{row}}$ $=\frac{1}{q\sum_{i=1}^{n}w(\operatorname{MMD}(\bm{X}_{i}))}\sum_{i=1}^{n}w(\operatorname{MMD}(\bm{X}_{i}))(\bm{X}_{i}-\tilde{\bm{\operatorname{M}}})\tilde{\bm{\Omega}}^{\operatorname{col}}(\bm{X}_{i}-\tilde{\bm{\operatorname{M}}})^{\prime}\text{, and}$ (20)
$\tilde{\bm{\Sigma}}^{\operatorname{col}}$ $=\frac{1}{p\sum_{i=1}^{n}w(\operatorname{MMD}(\bm{X}_{i}))}\sum_{i=1}^{n}w(\operatorname{MMD}(\bm{X}_{i}))(\bm{X}_{i}-\tilde{\bm{\operatorname{M}}})^{\prime}\tilde{\bm{\Omega}}^{\operatorname{row}}(\bm{X}_{i}-\tilde{\bm{\operatorname{M}}}),$ (21)
where $w:[0,\infty)\rightarrow[0,\infty)$ is a non-increasing and bounded weight function such that $w(\operatorname{MMD}(\bm{X}_{i}))>0$ for at least $\lfloor\nicefrac{{(n+d+2)}}{{2}}\rfloor$ observations that vanishes for large distances, i.e., $w(\operatorname{MMD}(\bm{X}_{i}))=0$ if $\operatorname{MMD}(\bm{X}_{i})>c_{1}>0$.
The following theorem shows that the reweighted MMCD estimator preserves the breakdown point of the original estimator. The simulations presented in Section 6 illustrate substantial improvements in the efficiency of the reweighed MMCD estimators. With increasing sample size, the finite sample efficiency exceeds $90\%$ across various selections of $p$ and $q$.
@@
$\varepsilon^{\ast}(\tilde{\bm{\operatorname{M}}}_{\bm{\mathfrak{X}}},\bm{\mathfrak{X}})\geq\frac{m}{n}\quad\text{and}\quad\varepsilon^{\ast}(\tilde{\bm{\Sigma}}^{\operatorname{row}}_{\bm{\mathfrak{X}}},\tilde{\bm{\Sigma}}^{\operatorname{col}}_{\bm{\mathfrak{X}}},\bm{\mathfrak{X}})\geq\frac{m}{n}.$
@@ 4.1 Adapting the C-step
Adapting the structure of the Fast-MCD algorithm to the matrix-variate setting leads to the development of the MMCD algorithm. This adaptation necessitates a modification in the covariance estimation during the C-step to derive suitable counterparts for computing the MMCD estimators. However, this process encounters a challenge due to the involvement of two covariance matrices, as depicted in Equations (11) and (12), both lacking closed-form solutions for their estimation. To address this issue, we incorporate the flip-flop algorithm introduced by Dutilleul, (1999) within the C-step. Consider a matrix-variate random sample $\bm{\mathfrak{X}}=(\bm{X}_{1},\dots,\bm{X}_{n})$, with $\bm{X}_{i}\in\mathbb{R}^{p\times q}$, and any $h$-subset $H_{\operatorname{old}}\subset\{1,\ldots,n\}$, with $\left|H_{\operatorname{old}}\right|=h>\lfloor\nicefrac{{p}}{{q}}+\nicefrac{{q}}{{p}}\rfloor+2$. First, the MLEs $(\hat{\bm{\operatorname{M}}}_{H_{\operatorname{old}}},\hat{\bm{\Sigma}}^{\operatorname{row}}_{H_{\operatorname{old}}},\hat{\bm{\Sigma}}^{\operatorname{col}}_{H_{\operatorname{old}}})$ are computed based on the observations in the subset $H_{\operatorname{old}}$ using the flip-flop algorithm, which is non-decreasing in likelihood. Next, compute the squared Mahalanobis distances $d^{2}_{i}(H_{\operatorname{old}}):=\operatorname{MMD}^{2}(\bm{X}_{i},\hat{\bm{\operatorname{M}}}_{H_{\operatorname{old}}},\hat{\bm{\Sigma}}^{\operatorname{row}}_{H_{\operatorname{old}}},\hat{\bm{\Sigma}}^{\operatorname{col}}_{H_{\operatorname{old}}})$ for all $i=1,\ldots,n$. In Proposition 2.0.1, we showed that $\sum_{i\in H_{\operatorname{old}}}d^{2}_{i}(H_{\operatorname{old}})=hpq$, hence only the terms of the log-likelihood function involving the determinants change in this step. To construct the new subset $H_{\operatorname{new}}$, sort the squared MMDs in ascending order, resulting in a permutation $\pi$ of $\{1,\ldots,n\}$ such that $d^{2}_{\pi(1)}(H_{\operatorname{old}})\leq\ldots\leq d^{2}_{\pi(n)}(H_{\operatorname{old}})$, and define a new $h$-subset $H_{\operatorname{new}}=\{\pi(1),\ldots,\pi(h)\}$. Since the estimators do not change in this step, the terms involving the determinant are constant, and by construction, the sum of the Mahalanobis distances either decreases or stays constant. Hence, the reordering is non-decreasing in likelihood. Finally, the estimators are updated using the flip-flop algorithm based on the observations in the subset $H_{\operatorname{new}}$, resulting in estimators $(\hat{\bm{\operatorname{M}}}_{H_{\operatorname{new}}},\hat{\bm{\Sigma}}^{\operatorname{row}}_{H_{\operatorname{new}}},\hat{\bm{\Sigma}}^{\operatorname{col}}_{H_{\operatorname{new}}})$, increasing the likelihood once more, and it follows that
@@ 4.2 The MMCD algorithm
The MMCD algorithm is a matrix-variate extension of the Fast-MCD procedure of Rousseeuw and Driessen, (1999), aiming to alleviate the C-steps dependence on the initial subset by using multiple initial subsets, iteratively conducting C-steps on each until convergence, and ultimately selecting the solution with the lowest determinant. While this explains the idea of the algorithm, there are more computational considerations and adjustments in the full MMCD algorithm. A pseudo-code of the MMCD Algorithm 2 is given in Supplement C.
@@
As in its multivariate counterpart, the MMCD procedure uses so-called elemental subsets to initialize the procedure. This means that we use $m$ subsets of size $d+2,d=\lfloor\nicefrac{{p}}{{q}}+\nicefrac{{q}}{{p}}\rfloor$, instead of size $h$, to increase the probability of obtaining at least one clean initial subset. Using $m=500$ elemental subsets by default allows for a reasonable tradeoff between a wide variety of settings where we likely obtain at least one clean elemental subset and the computational demands of computing initial estimators. If either $p\ll q$ or $q\gg p$, $d$ will be large, and using more initial subsets is recommended. Using elemental subsets increases not only the robustness of the initial estimators but also the computational efficiency.
@@
Moreover, the MMCD procedure only uses 2 C-step and MLE iterations for the initial elemental subsets to ensure even faster computation of the initial estimators. In the MLE procedure, Werner et al., (2008) demonstrated that the same asymptotic efficiency can be attained using only two iterations instead of iterating until convergence. As for the C-step, Rousseeuw and Driessen, (1999) outlined that after two iterations, subsets with the lowest covariance determinant during the procedure can already be identified, even before reaching convergence. Moreover, simulations show that we can identify those initial subsets that yield robust solutions after 2 C-steps, whether we use 2 MLE iterations or iterate the flip-flop algorithm until convergence. This is described in detail in Supplement C.1, where we also show that elemental subsets indeed yield more robust solutions than their larger counterparts in case of high contamination.
@@
The initialization step of the MMCD procedure yields $m$ initial estimators, and we keep the $10$ estimators with the lowest covariance determinant. Using those as initial estimators, we iterate C-steps until convergence on the complete data set $\bm{\mathfrak{X}}$. The solution with the lowest covariance determinant then yields the raw MMCD estimators.
@@
The raw MMCD estimators are scaled using the consistency factor $c(\alpha)$ given in Equation (18) to achieve consistency at the normal model as outlined in Theorem 3.0.2. Based on those rescaled raw MMCD estimators, the reweighted estimators described in Definition 3.0.1 are computed using the weights given in Equation (22). The reweighted MMCD estimators are then scaled using $c(\tilde{\alpha})=c(\nicefrac{{\tilde{h}}}{{n}})$, where $\tilde{h}$ denotes the number of observations with weights one.
@@
The MMCD algorithm repeatedly computes Mahalanbois distances for all $n$ samples, which is computationally expensive when $n$ gets large. To improve the computational efficiency for settings where $n$ is large, we implemented the subsampling approach proposed by Rousseeuw and Driessen, (1999). The idea is to split the sample of $n$ observations into several smaller subsamples and compute initial estimators on those subsamples before working on the large set with $n$ observations.